# Automated Modeling Method for Pathloss Model Discovery

### v2

@@ III-B The Indoor and Outdoor Empirical Models
El Chall et al.[35] derived empirical PL models for LoRaWAN (Long-Range Wide Area Network) based on extensive field measurements conducted in Lebanon. Indoor experiments were carried out at Saint Joseph University, while outdoor experiments were performed in Beirut City. The collected data were analyzed using linear regression to derive PL models for both indoor and outdoor environments.
@@
where $L_{h}$ denotes the additional loss related to the antenna height $h_{\text{ED}}$. The parameters obtained from the fitting process are: PL exponent $n=3.119$, reference PL $PL_{0}=140.7$, and antenna height loss factor $L_{h}=-4.7$, $X_{\sigma}=9.7$ is a zero-mean Gaussian random variable representing shadow fading.
@@ IV Methodologies
The design decision to adopt KANs as candidate techniques for enabling the automation of PL learning is justified by the shortcomings of typical techniques in learning logarithmic functions [5]. Knowing that existing PL models often involve logarithmic functions in their expression [8], KANs may have superior performance in approximating such models while being able to produce two flavors of interpretability.
@@ IV-A Kolmogorov-Arnold Networks
KANs have recently emerged as an alternative to multi-layer perceptron (MLP) based architectures, demonstrating strong performance on scientific tasks while maintaining interpretability. KANs have strong mathematical foundations based on the Kolmogorov-Arnold representation theorem, which states that any multivariate continuous function can be expressed as a sum of univariate functions [27]. KANs and MLPs implement activation functions in fundamentally different ways: KANs have activation functions on edges, while MLPs have activation functions on nodes. This unique design allows KANs to provide a more transparent and interpretable mapping between inputs and outputs, enhancing explainability while preserving the model’s ability to capture complex patterns.
The structure of a KAN can be represented as $[n_{1},…,n_{L+1}]$, where $L$ denotes the total number of layers. A deeper KAN is constructed by composing $L$ layers as follows:
$Y=\text{KAN}(X)=(\Phi_{L}\circ\Phi_{L-1}\circ\dots\circ\Phi_{1})X.$ (8)
@@
$\Phi=\{\phi_{q,p}\},\;p=1,2,\dots,d_{in},\;q=1,2,\dots,d_{out}.$ (9)
KAN utilizes the SiLU activation function combined with B-splines to enhance expressiveness. In this setup, edges control the transformations between layers, while nodes perform simple summation operations. A B-spline of order $k$ requires $G+k$ basis functions to define the spline over a given grid. Consequently, for each input (node in a layer), evaluating a B-spline of order $k$ involves computing $G+k-1$ basis functions and performing a weighted sum with the corresponding control points:
$\text{spline}(x)=\sum_{i=0}^{G+k-1}c_{i}B_{i}(x).$ (10)
@@ IV-B Deep Symbolic Regression
At the core of DSR [7] are the representation of mathematical expressions as sequences, the development of an autoregressive model to generate these expressions, and the use of a risk-seeking policy gradient approach to train the model to produce more accurate and well-fitting expressions.
@@
$p(\tau_{i}\mid\tau_{j\neq i},\theta)=p(\tau_{i}\mid\tau_{j<i},\theta).$ (11)
The sequence generator is generally trained using RL or related techniques. From this perspective, the sequence generator works as a RL policy to be optimized. This process involves sampling a batch of $N$ expressions $\mathcal{T}$, evaluating each expression using a reward function $R(\tau)$, and applying gradient descent to minimize a loss function. In this work, we employ three approaches to train the RNN.
@@ IV-B1 Risk-Seeking Policy Gradient (RSPG)
Originally proposed in [7], this approach prioritizes optimizing the best-case reward rather than the average. The loss function is defined as:
@@ IV-B2 Vanilla Policy Gradient (VPG)
This method employs the REINFORCE algorithm [36], where training is performed over the batch $\mathcal{T}$ with the loss function:
@@ IV-B3 Priority Queue Training (PQT)
Proposed by Abolafia [37], this non-RL approach focuses on optimizing best-case performance. It stores samples from each batch in a maximum reward priority queue (MRPQ), and training is conducted on these stored samples using a supervised learning objective:
@@
$R(\tau)=\frac{1}{1+\text{NRMSE}(\tau)}.$ (16)
@@ V Evaluation Methodologies
To ensure full replicability of the entire study and elaborate on the setup under which the evaluation has been done, in this section, we first discuss the type of data used to train the automatic discovery methods, followed by the training methodologies for each method.
@@ V-A1 Synthetic Data
The dataset was generated following the approach described in [32] to approximate the PL models defined by Eq. (1) and Eq. (2) in Section III-A. The parameter ranges, summarized in Table I, were derived from authentic data in studies that comprehensively analyze various scenarios and frequency ranges, highlighting the applicability of $PL^{ABG}$ and $PL^{CI}$ models in various contexts [9, 34, 38]. Specifically, for $PL^{ABG}$ the parameters $\alpha$,$\gamma$, $\beta$, $f$, $d$, and $\chi$ were used as inputs, while for $PL^{CI}$ $f$,$n$, $d$, and $\chi$ were the inputs. The dataset consists of 1,000 instances. The frequency was converted to hertz for the $PL^{CI}$, following standard practice in the literature.
@@
$\boldsymbol{\alpha}$ $[0.1,2.5]$
$\boldsymbol{\beta}$ $[-10,-1]$
$\boldsymbol{\gamma}$ $[0,2]$
@@ V-A2 Empirically Measured Data
Measurements were conducted in two distinct environments, indoor and outdoor, as described in [35]. The indoor dataset includes 1,317 instances, and the outdoor dataset contains 787 instances. Input parameters for $PL^{EI}$ included $n_{w}$, $n_{f}$, $d$, and $f$ , while the $PL^{EO}$ used $h_{\mathrm{ED}}$, $d$, and $f$. The corresponding parameter ranges are presented in Table II. For performance evaluation, we compared the automated methods with the models described in Section III-B and developed by the authors in [35] using the real-world measurements [^2] https://zenodo.org/records/1560654.
@@
$\boldsymbol{h}_{\text{ED}}(m)$ $[0.2,3]$
Figure 2: The flow diagram of developing a pathloss model using KANs.
Figure 3: The flow diagram of developing a pathloss model using DSR.
@@ V-B Training of KANs
We use the original KAN implementation from the pykan library [^3] https://github.com/KindXiaoming/pykan and train the KANs in a similar way as other ML models, using the 80-20 split of data as depicted in Figure 2. During training, various configurations of layers and grid sizes are varied to optimize performance. In this process, the splines and weights of the edges connecting these splines, as described in Section IV-A, are learned. Unlike MLPs, KANs are not very deep (i.e., usually about 3 layers with 5 being very deep) and not very wide [27]. Due to their relative shallowness and narrowness, once learned, they can be visualized as a graph with the learned activation functions overlayed on the edges and the importance of compositions represented as edge weight. This alone already gives more intuition on the learned model than in the case of other neural architectures. However, as a second step, a mechanism to match the learned splines with actual known functions also exists. We perform default automated matching relying on the library’s default $auto\_symbolic()$ functionality. After the matching step, the network is retrained with the respective functions by only updating the weights of the edges. Finally, a fully interpretable symbolic formula is extracted as a last step in the training. To tune the KANs, we perform a full grid search for each model with grid size $grids\in\{5,8,10,15,20,30,40,50\}$, $steps\in\{50,100,200,300\}$ and $lambda\in\{0.02,0.002,0.0002,0.00002\}$. The final tuned hyperparameters for the PL models learned by KANs are summarized in Table III. For both the CI and ABG models, we normalize the training data to maintain consistency with our previous work [32]. In preliminary experiments without normalization, the frequency, expressed in hertz in the CI model, tended to dominate the other input variables, resulting in imbalanced feature contributions.
@@
Layers Grid Steps k $\boldsymbol{\lambda}$
@@ V-C Training of DSR
In DSR, a set of constraints can be implemented to narrow the search space, ensuring the generation of meaningful expressions while preserving computational efficiency. The following constraints were specifically applied. Expression Length Limits: Expressions were constrained to a minimum length of 4 to avoid trivial solutions and a maximum length of 40 to maintain interpretability. Avoiding Redundant Constants: Operators were restricted to prevent all their children from being constants, preventing the generation of expressions that simplify to a single constant. Unary Operator Consistency: Unary operators were restricted from having their inverse as a child (e.g., $\log(\exp(x))$) to prevent the generation of redundant expressions. Trigonometric Operator Composition: Trigonometric operators were restricted from having other trigonometric operators as descendants (e.g., $\sin(1+\cos(x))$) to prevent unnecessary complexity, despite being mathematically valid. Repeat Constraint: This constraint regulates the occurrence of specific tokens (e.g., operators and/or variables) in generated expressions by enforcing minimum and maximum limits. It softly penalizes violations of the minimum threshold to allow exploration while strictly invalidating expressions that exceed the maximum limit. The repeat constraint with $f$ and $d$ was utilized in all cases.
Each of the three RNN-based algorithms (RSPG, PQT, and VPG) has one unique algorithm-specific parameter, for which we adopted default values as presented in [7], i.e. the risk factor ($\epsilon=0.05$) for RSPG, the priority queue size ($k=10$) for PQT, and the EWMA coefficient ($\alpha=0.25$) for VPG. In addition to these unique parameters, all three algorithms share a common set of hyperparameters (batch size, learning rate and entropy weight). We explored various values for these hyperparameters to achieve the best performance. The final tuned hyperparameters are summarized in Table IV.
@@ VI-A Learning the ABG Model from Synthetic Data
The performance comparison for KAN and DSR methods on the $PL^{ABG}$ is presented in Table V. The KAN [6,6,1] achieves superior predictive performance, with exceptionally low error metrics (MAE = 0.06, MSE = 0.01, MAPE = 0.28) and $R^{2}$ value of 1, indicating a nearly perfect alignment between the model predictions and the actual data as confirmed by the plot in Figure 4. In the figure, pred stands for the predicted PL values and true stands for the actual values based on Eq. (1). In contrast, the KAN [6,6,1] Auto-symbolic, which replaces the learned splines with known functions and re-learns new weights for the edges of the graph, shows significantly lower accuracy, with MAE = 3.89, MSE = 28.14, MAPE = 12.51, and $R^{2}$ = 0.93. Among DSR models, DSR-RSPG obtained the best performance (MAE = 3.86, MSE = 33.50, MAPE = 14.28, $R^{2}$ = 0.92), followed by DSR-PQT (MAE = 5.90, MSE = 62.07, MAPE = 21.83, $R^{2}$ = 0.90) and DSR-VPG (MAE = 6.67, MSE = 83.38, MAPE = 20.48, $R^{2}$ = 0.77). These results highlight that while KAN [6,6,1] achieves the highest accuracy, other methods show varying performance metrics, generally with higher error rates and lower values of $R^{2}$.
@@
(a) Train loss (the smaller, the better).
(b) Test loss (the smaller, the better).
Figure 6: Parameter search for the KAN [6,6,1] architecture for approximating $PL^{ABG}$.
@@
Methods Metrics MAE MSE MAPE R2 Expression
KAN [6,6,1] 0.06 0.01 0.28 1 Tree and splines in Fig. 5
@@
KAN [6,6,1] Auto-symbolic 3.89 28.14 12.51 0.93 $-291.519\operatorname{log}_{10}{\left(8.686-3.052\alpha\right)}+32.175\cos{\left(1.116\gamma-1.323\right)}+9.005\beta-53.258\operatorname{log}_{10}{\left(9.367-3.859d\right)}+0.016\left(0.028f-1\right)^{2}+313.289-\frac{0.476}{\left(-f-0.168\right)^{2}}+8.008\chi$
DSR-RSPG 3.86 33.50 14.28 0.92 $20\alpha+10\gamma+\beta+\log_{10}(d)+\frac{\gamma f}{10}+\chi$
DSR-PQT 5.90 62.07 21.83 0.90 $20\alpha+15\gamma+\beta+\frac{f}{20}+\chi$
DSR-VPG 6.67 83.38 20.48 0.77 $(\alpha+\gamma)(10+\chi)$
@@
As discussed in Sections IV-A and V-B, KANs learn splines and weights on the edges, and the best configuration is found using grid search. In our experiments, the best KAN architecture learned for $PL^{ABG}$ is depicted in Figure 5 and has a three-layer design [6, 6, 1]. The first layer includes six input nodes, corresponding to the parameters $\alpha,\gamma,\beta,f,d,\chi$; the second layer has the same number of nodes, and the output layer consists of a single node representing the predicted value of the function. This architecture was selected based on the smallest train losses and test losses during the grid search to find this model. The result of the grid search is illustrated in Figure 6 (a and b), showing how variations in these parameters affect test loss and train loss, with color gradients illustrating the scale of the losses.
@@
It can be seen from the learned graph in Figure 5 that the final expression is the summation of four terms, with the third one having a lower weight, i.e. very light gray coloring of the edge that connects the hidden layer with the output layer. It can also be observed from the functions learned on the first three edges of the last layer of the graph that they perform a non-linear transformation, while the last one performs a linear one. From the coloring of the edges and summations in the hidden layer, the first summation node, a quasi-linear transformation of $\alpha$ is summed with a non-linear transformation of $d$ with medium weighting. In the second summation node, a medium weighted quasi-linear transformation of $\gamma$ with a non-linear transformation of $f$. The next three summation nodes have low to minimal weighting, while the last summation node aggregates all six inputs with high to medium weights.
@@
Finally, we examine the validity of the expressions generated by the KANs and DSR methods. For $PL^{ABG}$, the expressions obtained by KAN [6,6,1] Auto-symbolic and DSR-RSPG are considered valid, as they correctly represent PL as a function of $f$ and $d$. We also notice that the expression learned by the KAN [6,6,1] Auto-symbolic is more complex than the one learned by the DSR-RSPG. In contrast, the expression generated by DSR-PQT fails to include $d$, and the DSR-VPG expression excludes both $f$ and $d$. These results indicate that the DSR-PQT and DSR-VPG expressions are unlikely to be valid for accurately modeling PL from a communication engineering perspective.
TABLE VI: Performance comparison on CI model.
Methods Metrics MAE MSE MAPE R2 Expression
KAN [4,4,1] 0.05 0.01 0.04 1 Tree and splines in Figure 8
KAN [4,4,1] Auto-symbolic 4.56 36.26 3.05 0.96 $-280.464\operatorname{log}_{10}{\left(8.729-1.621n\right)}+43.806\operatorname{log}_{10}{\left(1.064d+1.913\right)}+19.929\operatorname{log}_{10}{\left(5.162f+9.329\right)}+50.758\operatorname{log}_{10}{\left(0.789\chi+7.457\right)}+352.246$
DSR-RSPG 11.49 261.46 10.21 0.80 $4023n+\frac{d}{10}+log_{10}(f)+$ 40
DSR-PQT 13.60 223.05 10.58 0.84 $4023n+\frac{d}{10}+(\chi+n)cos(log_{10}(f))+$ 40
DSR-VPG 16.90 570.92 15.24 0.67 $6021n+2\chi+$ 60

### v3

@@ III-B The Indoor and Outdoor Empirical Models
El Chall et al.[ElChall] derived empirical PL models for LoRaWAN (Long-Range Wide Area Network) based on extensive field measurements conducted in Lebanon. Indoor experiments were carried out at Saint Joseph University, while outdoor experiments were performed in Beirut City. The collected data were analyzed using linear regression to derive PL models for both indoor and outdoor environments.
@@
where $L_{h}$ denotes the additional loss related to the antenna height $h_{\text{ED}}$. The parameters obtained from the fitting process are: PL exponent $n=3.119$, the reference PL $PL_{0}=140.7$, and the antenna height loss factor $L_{h}=-4.7$, $X_{\sigma}$ is a variable that represents shadow fading which takes a random value following the Gaussian distribution with zero mean and standard deviation of $\sigma$, where $\sigma$=9.7.
@@ IV Methodologies
The design decision to adopt KANs and DSR as candidate techniques to enable the automation of PL learning is justified by the shortcomings of typical techniques in learning logarithmic functions [makke2024interpretable]. Knowing that existing PL models often involve logarithmic functions in their expression [Phillips], KANs and DSR may have superior performance in approximating such models while being able to produce different flavors of interpretability.
Overall, different methods provide varying levels of interpretability. DNNs offer low interpretability through their activation maps. In contrast, KANs offer a medium level of interpretability through their spline-based representations. DSR and KANs (by using Auto-symbolic mapping) can achieve high interpretability by expressing models as explicit mathematical functions.
@@ IV-A Kolmogorov-Arnold Networks
KANs have recently emerged as an alternative to multi-layer perceptron (MLP) based architectures, demonstrating strong performance on scientific tasks while maintaining interpretability. KANs have strong mathematical foundations based on the Kolmogorov-Arnold representation theorem, which states that any multivariate continuous function can be expressed as a sum of univariate functions [Liukan]. KANs and MLPs implement activation functions in fundamentally different ways: KANs have activation functions on edges, while MLPs have activation functions on nodes. This unique design allows KANs to provide a more transparent and interpretable mapping between inputs and outputs, enhancing explainability while preserving the model’s ability to capture complex patterns.
The structure of a KAN can be represented as $n_{1},…,n_{L+1}]$, where $L$ denotes the total number of layers. A deeper KAN is constructed by composing $L$ layers as follows:
$Y=\text{KAN}(X)=(\Phi_{L}\circ\Phi_{L-1}\circ\dots\circ\Phi_{1})X,$ (8)
@@
$\Phi=\{\phi_{q,p}\},\;p=1,2,\dots,d_{in},\;q=1,2,\dots,d_{out},$ (9)
KAN utilizes the SiLU activation function combined with B-splines to enhance expressiveness. In this setup, edges control the transformations between layers, while nodes perform simple summation operations. A B-spline of order $k$ requires $G+k$ basis functions to define the spline over a given grid. Consequently, for each input (node in a layer), evaluating a B-spline of order $k$ involves computing G+k1G+k-1 basis functions and performing a weighted sum with the corresponding control points:
$\text{spline}(x)=\sum_{i=0}^{G+k-1}c_{i}B_{i}(x),$ (10)
@@ IV-B Deep Symbolic Regression
At the core of DSR [Petersen] are the representation of mathematical expressions as sequences, the development of an autoregressive model to generate these expressions, and the use of a risk-seeking policy gradient approach to train the model to produce more accurate and well-fitting expressions.
@@
$p(\tau_{i}\mid\tau_{j\neq i},\theta)=p(\tau_{i}\mid\tau_{j<i},\theta),$ (11)
The sequence generator is generally trained using RL or related techniques. From this perspective, the sequence generator works as an RL policy to be optimized. This process involves sampling a batch of $N$ expressions $\mathcal{T}$, evaluating each expression using a reward function $R(\tau)$, and applying gradient descent to minimize a loss function. In this work, we employ three approaches to train the RNN.
@@ IV-B1 Risk-Seeking Policy Gradient (RSPG)
Originally proposed in [Petersen], this approach prioritizes optimizing the best-case reward rather than the average. The loss function is defined as:
@@ IV-B2 Vanilla Policy Gradient (VPG)
This method employs the REINFORCE algorithm [Williams], where training is performed over the batch $\mathcal{T}$ with the loss function:
@@ IV-B3 Priority Queue Training (PQT)
Proposed by Abolafia [Abolafia], this non-RL approach focuses on optimizing best-case performance. It stores samples from each batch in a maximum reward priority queue (MRPQ), and training is conducted on these stored samples using a supervised learning objective:
@@
$R(\tau)=\frac{1}{1+\text{NRMSE}(\tau)},$ (16)
@@ V Evaluation Methodologies
To ensure complete replicability of the study and clarify the conditions under which the evaluation was conducted, this section first discusses the type of data used to train the automatic discovery methods, followed by the training methodologies for each method.
@@ V-A1 Synthetic Data
The dataset was generated following the approach described in [anaqreh] to approximate the PL models defined by Eq. (1) and Eq. (2) in Section III-A. The parameter ranges, summarized in Table I, were derived from authentic data in studies that comprehensively analyze various scenarios and frequency ranges, highlighting the applicability of $PL^{ABG}$ and $PL^{CI}$ models in various contexts [Sun, Sun2, Cheng]. Specifically, for $PL^{ABG}$ the parameters $\alpha$$\gamma$, $\beta$, $f$, $d$, and $\chi^{ABG}_{\sigma}$ were used as inputs, while for $PL^{CI}$ $f$,$n$, $d$, and $\chi^{CI}_{\sigma}$ were the inputs. The dataset consists of 1,000 instances. The frequency was converted to hertz for the $PL^{CI}$, following standard practice in the literature.
@@
$\boldsymbol{\alpha}$ $[0.1,2.5]$ -
$\boldsymbol{\beta}$ [10,1][-10,-1] -
$\boldsymbol{\gamma}$ $[0,2]$ -
@@ V-A2 Empirically Measured Data
El Chall et al. [ElChall] conducted measurements in both indoor and outdoor environments. The indoor dataset includes 1,317 instances, and the outdoor dataset contains 787 instances. Input parameters for $PL^{EI}$ included $n_{w}$, $n_{f}$, $d$, and $f$ , while the $PL^{EO}$ used $h_{\mathrm{ED}}$, $d$, and $f$. The corresponding parameter ranges are presented in Table II. For performance evaluation, we compared the automated methods with the models described in Section III-B and developed by the authors in [ElChall] using the real-world measurements [^2] https://zenodo.org/records/1560654.
@@
$\boldsymbol{h}_{\text{ED}}(m)$ - $[0.2,3]$
@@ V-A3 Analytical Baselines
Figure 2: The flow diagram of developing a pathloss model using KANs.
@@ V-B Training of KANs
We used the original KAN implementation from the pykan library [^3] https://github.com/KindXiaoming/pykan and trained the KANs in a similar way as other ML models, using the 80-20 split of data as depicted in Figure 2. During training, layers and grid sizes are configured to optimize performance. In this process, the splines and weights of the edges connecting these splines, as described in Section IV-A, are learned. Unlike MLPs, KANs are not very deep (i.e., usually about 3 layers with 5 being very deep) and not very wide [Liukan]. Due to their relative shallowness and narrowness, once learned, they can be visualized as a graph with the learned activation functions overlayed on the edges and the importance of compositions represented as edge weight. This alone already gives more intuition on the learned model than in the case of other neural architectures. However, as a second step, a mechanism to match the learned splines with actual known functions also exists. We perform default automated matching relying on the library’s default auto_symbolic()auto\_symbolic() functionality. After the matching step, the network is retrained with the respective functions by only updating the weights of the edges. Finally, a fully interpretable symbolic formula is extracted as a last step in the training. To tune the KANs, we perform a full grid search for each model with $grids\in\{5,8,10,15,20,30,40,50\}$, $steps\in\{50,100,200,300\}$ and $lambda\in\{0.02,0.002,0.0002,0.00002\}$.
@@
Layers Grids Steps k $\boldsymbol{\lambda}$
@@
The final tuned hyperparameters for the PL models learned by KANs are summarized in Table III. For both the CI and ABG models, we normalize the training data to maintain consistency with our previous work [anaqreh]. In preliminary experiments without normalization, the frequency, expressed in hertz in the CI model, tended to dominate the other input variables, resulting in imbalanced feature contributions.
Figure 3: The flow diagram of developing a pathloss model using DSR.
@@ V-C Training of DSR
In DSR, a set of constraints can be implemented to narrow the search space, ensuring the generation of meaningful expressions while preserving computational efficiency. The following constraints were specifically applied. Expression Length Limits: Expressions were constrained to a minimum length of 4 tokens to avoid trivial solutions and a maximum length of 40 tokens to maintain interpretability. Avoiding Redundant Constants: Operators were restricted to prevent all their children from being constants, preventing the generation of expressions that simplify to a single constant. Unary Operator Consistency: Unary operators were restricted from having their inverse as a child (e.g., $\log(\exp(x))$) to prevent the generation of redundant expressions. Trigonometric Operator Composition: Trigonometric operators were restricted from having other trigonometric operators as descendants (e.g., $\sin(1+\cos(x))$) to prevent unnecessary complexity, despite being mathematically valid. Repeat Constraint: This constraint regulates the occurrence of specific tokens (e.g., operators and/or variables) in generated expressions by enforcing minimum and maximum limits. It softly penalizes violations of the minimum threshold to allow exploration while strictly invalidating expressions that exceed the maximum limit. The repeat constraint with $f$ and $d$ was utilized in all cases.
Each of the three RNN-based algorithms (RSPG, PQT, and VPG) has one unique algorithm-specific parameter, for which we adopted default values as presented in [Petersen], i.e., the risk factor ($\epsilon=0.05$) for RSPG, the priority queue size ($k=10$) for PQT, and the EWMA coefficient ($\alpha=0.25$) for VPG. In addition to these unique parameters, all three algorithms share a common set of hyperparameters (batch size, learning rate, and entropy weight). We explored various values for these hyperparameters to achieve the best performance. The final tuned hyperparameters are summarized in Table IV.
@@ VI-A Learning the ABG Model from Synthetic Data
The performance comparison for the KAN and DSR methods on $PL^{ABG}$ is presented in Table V. The KAN [6,6,1] achieves superior predictive performance, with exceptionally low error metrics (MAE = 0.06, MSE = 0.01, MAPE = 0.28) and $R^{2}$ value of 1, indicating a nearly perfect alignment between the model predictions and the actual data as confirmed by the plot in Figure 4. In the figure, pred stands for the predicted PL values and true stands for the actual values based on Eq. (1). In contrast, the KAN [6,6,1] Auto-symbolic, which replaces the learned splines with known functions and re-learns new weights for the edges of the graph, shows significantly lower accuracy, with MAE = 3.89, MSE = 28.14, MAPE = 12.51, and $R^{2}$ = 0.93. Among DSR models, DSR-RSPG obtained the best performance (MAE = 3.86, MSE = 33.50, MAPE = 14.28, $R^{2}$ = 0.92), followed by DSR-PQT (MAE = 5.90, MSE = 62.07, MAPE = 21.83, $R^{2}$ = 0.90) and DSR-VPG (MAE = 6.67, MSE = 83.38, MAPE = 20.48, $R^{2}$ = 0.77). These results highlight that while KAN [6,6,1] achieves the highest accuracy, other methods show varying performance metrics, generally with higher error rates and lower values of $R^{2}$.
@@
Methods Metrics MAE MSE MAPE R2 Expression Interpretability Validity
KAN [6,6,1]* 0.06 0.01 0.28 1 Tree and splines in Fig. 5 Medium ?
@@
KAN [6,6,1] Auto-symbolic* 3.89 28.14 12.51 0.93 $-291.5\operatorname{log}_{10}{\left(8.7-3.1\alpha\right)}+32.2\cos{\left(1.1\gamma-1.3\right)}+9\beta-53.3\operatorname{log}_{10}{\left(9.4-3.9d\right)}+0.02\left(0.03f-1\right)^{2}+313.3-\frac{0.5}{\left(-f-0.2\right)^{2}}+8\chi_{\sigma}$ High
DSR-RSPG 3.86 33.50 14.28 0.92 $20\alpha+10\gamma+\beta+\log_{10}(d)+\frac{\gamma f}{10}+\chi_{\sigma}$ High
DSR-PQT 5.90 62.07 21.83 0.90 $20\alpha+15\gamma+\beta+\frac{f}{20}+\chi_{\sigma}$ High
DSR-VPG 6.67 83.38 20.48 0.77 $(\alpha+\gamma)(10+\chi_{\sigma})$ High
* Input parameters were normalized by dividing each value by the maximum value of its corresponding parameter
Interpretability
@@
As discussed in Sections IV-A and V-B, KANs learn splines and weights on the edges, and the best configuration is found using grid search. In our experiments, the best KAN architecture learned for $PL^{ABG}$ is depicted in Figure 5 and has a three-layer design [6,6,1]. The first layer includes six input nodes, corresponding to the parameters $\alpha,\gamma,\beta,f,d,\chi_{\sigma}$; the second layer has the same number of nodes, and the output layer consists of a single node representing the predicted value of the function.
@@
It can be seen from the learned graph in Figure 5 that the final expression is the summation of four terms, with the third one having a lower weight, i.e. very light gray coloring of the edge that connects the hidden layer with the output layer. The fourth and fifth edges from the last layer were learned as being unimportant, resulting in the disconnection of their associated summation nodes from the overall composition graph. It can also be observed from the functions learned on the first three edges of the last layer of the graph that they perform a nonlinear transformation, while the last one performs a linear one. From the coloring of the edges and summations in the hidden layer, the first summation node, a quasi-linear transformation of $\alpha$ is summed with a nonlinear transformation of $d$ with medium weighting. In the second summation node, a medium weighted quasi-linear transformation of $\gamma$ with a nonlinear transformation of $f$. The next three summation nodes have low to minimal weighting, while the last summation node aggregates all six inputs with high to medium weights.
Validity
@@
We examine the validity of the expressions generated by the KANs and DSR methods. For $PL^{ABG}$, the expressions obtained by KAN [6,6,1] Auto-symbolic and DSR-RSPG are considered valid, as they correctly represent PL as a function of $f$ and $d$. We also notice that the expression learned by the KAN [6,6,1] Auto-symbolic is more complex than the one learned by the DSR-RSPG. Further methodological tools are needed as future work to assess the validity of the composition graph and spline from 5. In contrast, the expression generated by DSR-PQT does not include $d$, and the expression of DSR-VPG excludes both $f$ and $d$. These results indicate that the DSR-PQT and DSR-VPG expressions are unlikely to be valid to accurately model PL from a communication engineering perspective.
VI-B Learning the CI Model from Synthetic Data